MainPage:Nuclear:KaonDetector:PMTCharacteristics:LogBook:pag01

From cua_phy
Jump to navigation Jump to search

⇐ Back to the PMT Characteristics page
⇒ Go to Summer 2012 work page

LogBook on the CODA data acquisition, at CUA

Page (last), [6], [5], [4], [3], [2], 1.

LogBook - Page 01
Run Number Date User PMT High Voltage Experiment Description and observations Image
455 02/20/2012 Mike / Marco 09641 -1.78 kV PMT Scanning Reference PMT inserted. Four layers of office paper inserted in front of the LED, inside the hole for light colimating. Flashing the LED in front of the center of the 5" PMT with 10Hz, all night long. LED pulses of 5V x 20ns. Orientation: high voltage cable in the direction of the entrance door. REF in ch#2 and 5"PMT in ch#3. PMT S/N 09641 and base #101 -
456 02/21/2012 Mike / Marco 09641 -1.78 kV PMT Scanning Scanning the PMT with horizontal and vertical scanning. 25 x 25 position matrix. 500 points in each position. Same configuration of run #455.
Error creating thumbnail: Unable to save thumbnail to destination
467 02/24/2012 Marco 09641 -1.78 kV PMT Scanning Shielding inserted, making two 25x25 scans. The first one runs the lines and the second one runs the columns. Orientation: high voltage cable close in the direction of the back wall.
Error creating thumbnail: Unable to save thumbnail to destination
468 02/24/2012 Marco 09641 -1.78 kV PMT Scanning Same as #467, but rotating PMT by 180 degrees. Orientation: high voltage cable in the direction of the entrance door.
Error creating thumbnail: Unable to save thumbnail to destination
469 02/26/2012 Mike 09641 -1.78 kV PMT Scanning HV cable toward entrance, PMT tilted downward by about 15 degrees. No shielding.
Error creating thumbnail: Unable to save thumbnail to destination
470 02/26/2012 Mike 09641 -1.78 kV PMT Scanning Same as #469, but with PMT tilted upward by about 15 degrees. No shielding.
Error creating thumbnail: Unable to save thumbnail to destination
472 02/27/2012 Mike / Marco 09650 -1.78 kV PMT Scanning Changing the PMT to see if we have the same shape of the scanning (to compare with run #468, for the previous PMT). HV cable toward entrance. Shielding inserted. PMT S/N 09650 and base #102
Error creating thumbnail: Unable to save thumbnail to destination
473 02/27/2012 Mike / Marco 09650 -1.91 kV PMT Scanning Same as #472, just changing the High Voltage
Error creating thumbnail: Unable to save thumbnail to destination
474 02/27/2012 Mike / Marco 09650 -1.91 kV PMT Scanning Rotating the PMT by 180^o. New orientation: HV cable toward back wall.
Error creating thumbnail: Unable to save thumbnail to destination
475 02/27/2012 Mike / Marco 09650 -1.78 kV PMT Scanning Same as #474, just changing the High Voltage. It is a 180 degrees rotation of the run #472.
Error creating thumbnail: Unable to save thumbnail to destination
476 03/12/2012 Mike / Marco 09650 -1.78 kV PMT Gain Looking for the Single Electron Peak, flashing the LED in front of the PMT. Configuration: no shielding, PMT not oriented to some direction, 5V x 20ns pulses, PMT S/N 09650 with base #102. Conclusion: SEP not defined. Do we need to amplify the signal to increase the resolution of our system? Decrease the light intensity?
Error creating thumbnail: Unable to save thumbnail to destination
477 03/13/2012 Marco 09650 -1.78 kV PMT Gain Same as #476, changing the LED pulse: 3V @ 6ns.
Error creating thumbnail: Unable to save thumbnail to destination
478 03/13/2012 Marco 09650 -1.78 kV PMT Gain Junk. Forgot to turn LED on. -
479 03/13/2012 Marco 09650 -1.78 kV PMT Gain Same as #476, changing the LED pulse: 4V @ 6ns. About 1M samples @ 1kHz. Distance from pedestal to SEP: 15.63 channels.
Error creating thumbnail: Unable to save thumbnail to destination
480 03/13/2012 Marco 09650 -2.02 kV PMT Gain Same as #479, changing the High Voltage.
Error creating thumbnail: Unable to save thumbnail to destination
481 03/13/2012 Marco 09650 -1.89 kV PMT Gain Same as #480, changing the High Voltage.
Error creating thumbnail: Unable to save thumbnail to destination
488 03/14/2012 Felipe / Marco - - Setup Attenuation Factor First tentative to register the pulse in the ADC (see description at #489). Signal much intense, ADC saturated. -
489 03/14/2012 Felipe / Marco - - Setup Attenuation Factor Negative pulses were generated using the Avtec module. They were passing through all the cables and connectors to enter into the same channel #3 of the ADC. So the attenuation factor would be calculated. Gate width: 40ns. Pulse width: 10ns (beginning 5ns after the gate edge). Pulse area: -1.430(3)nVs. Pulse peak-to-peak: 133.6mV. Oscilloscope screen in 120314_153220.bmp. See 120314_153041.bmp used in the adjust of the gate. File:489.jpg
490 03/14/2012 Felipe / Marco - - Setup Attenuation Factor A jumper was inserted in the ADC channel #3 to measure the channel relative to area zero. The gate features remain the same as in #489.
Error creating thumbnail: Unable to save thumbnail to destination
491 03/14/2012 Felipe / Marco - - Setup Attenuation Factor Same as #489, changing the input pulses intensity. Pulse area: -826.9(30)pVs. Pulse peak-to-peak: 54.4mV. The width changed even without changing the pulse generator parameter: 14ns. Oscilloscope screen in 120314_153609.bmp.
Error creating thumbnail: Unable to save thumbnail to destination
492 03/14/2012 Felipe / Marco - - Setup Attenuation Factor Same as #489, changing the input pulses intensity. Pulse area: 515(10)pVs. Pulse peak-to-peak: 12.4mV. Pulse width: 38ns. Oscilloscope screen in 120314_171907.bmp.
Error creating thumbnail: Unable to save thumbnail to destination
493 03/14/2012 Felipe / Marco - - Setup Attenuation Factor Same as #489, changing the input pulses intensity. Pulse area: -2.876(20)nVs. Pulse peak-to-peak: 322mV. Pulse width: 7.80ns. Oscilloscope screen in 120314_172349.bmp.
Error creating thumbnail: Unable to save thumbnail to destination
494 03/14/2012 Felipe / Marco - - Setup Attenuation Factor Same as #489, changing the input pulses intensity. Pulse area: -4.290(20)nVs. Pulse peak-to-peak: 424mV. Pulse width: 8.16ns. Oscilloscope screen in 120314_172626.bmp.
Error creating thumbnail: Unable to save thumbnail to destination
495 03/14/2012 Felipe / Marco - - Setup Attenuation Factor Same as #489, changing the input pulses intensity. Pulse area: -5.630(20)nVs. Pulse peak-to-peak: 496mV. Pulse width: 9.58ns. Oscilloscope screen in 120314_172809.bmp.
Error creating thumbnail: Unable to save thumbnail to destination
496 03/14/2012 Felipe / Marco - - Setup Attenuation Factor Same as #489, changing the input pulses intensity. Pulse area: -1.064(10)nVs. Pulse peak-to-peak: 81.6mV. Pulse width: 10.95ns. Oscilloscope screen in 120314_173006.bmp.
Error creating thumbnail: Unable to save thumbnail to destination
497 03/14/2012 Felipe / Marco - - Setup Attenuation Factor Same as #496, changing the pulse and the gate widths to see if the integration is independent of the time. Pulse area: -1.180(20)nVs. Pulse width: n/a. Gate adjustment in 120314_173511.bmp.
Error creating thumbnail: Unable to save thumbnail to destination
- 03/14/2012 Felipe / Marco - - Setup Attenuation Factor Fitting for adjusting the attenuation factor. Using runs from #490 to #497. Adjusted attenuation factor: 0.9129 +/- 0.0220. Used factors: oscilloscope impedance 50 Ohm; ADC full range 256pC@10bits.
Error creating thumbnail: Unable to save thumbnail to destination
501 04/--/2012 Felipe 09650 -1.85 PMT Gain Flashing LED in the center of the PMT to analyze the gain.
Error creating thumbnail: Unable to save thumbnail to destination
508 04/--/2012 Felipe 09650 -1.95 PMT Gain Flashing LED in the center of the PMT to analyze the gain.
Error creating thumbnail: Unable to save thumbnail to destination
510 04/--/2012 Felipe 09650 -2.00 PMT Gain Flashing LED in the center of the PMT to analyze the gain.
Error creating thumbnail: Unable to save thumbnail to destination
512 04/--/2012 Felipe 09650 -2.05 PMT Gain Flashing LED in the center of the PMT to analyze the gain.
Error creating thumbnail: Unable to save thumbnail to destination
521 04/18/2012 Marco 09650 -1.79 PMT Scanning Scanning PMT with high resolution: 100x100 matrix, 10k points in each position. Time between flashes: 2ms (P-2). Grid size: 457 steps per movement. (lost 99 points in CODA, don't know why) -
522 04/19/2012 Marco 09650 -2.17 PMT Scanning Same as 521, just changing High Voltage. (lost 93 points) -
523 04/19/2012 Marco 09650 -2.30 PMT Scanning Same as 521, just changing High Voltage. -
524 04/20/2012 Marco 09650 -2.30 PMT Gain LED centered in front of the PMT flashing with 3V @ 10ns, rate approx. 1.1kHz. Configurations of the pulse generator module (Avtech AV-1030-C): fine pot amplitude in 7.0 and fine pot time width in 3.5; coarse pots in minumum.
Error creating thumbnail: Unable to save thumbnail to destination
525 04/20/2012 Marco 09650 -2.43 PMT Gain Same as 524, just changing High Voltage.
Error creating thumbnail: Unable to save thumbnail to destination
526 04/20/2012 Marco 09650 -1.67 PMT Gain Same as 524, just changing High Voltage.
Error creating thumbnail: Unable to save thumbnail to destination
527 04/20/2012 Marco 09650 -1.40 PMT Gain Same as 524, just changing High Voltage.
Error creating thumbnail: Unable to save thumbnail to destination
528 04/20/2012 Marco 09650 -1.50 PMT Gain Same as 524, just changing High Voltage.
Error creating thumbnail: Unable to save thumbnail to destination
529 04/20/2012 Marco 09650 -1.60 PMT Gain Same as 524, just changing High Voltage.
Error creating thumbnail: Unable to save thumbnail to destination
530 04/20/2012 Marco 09650 -1.70 PMT Gain Same as 524, just changing High Voltage.
Error creating thumbnail: Unable to save thumbnail to destination
531 04/20/2012 Marco 09650 -2.00 PMT Gain Same as 524, just changing High Voltage.
Error creating thumbnail: Unable to save thumbnail to destination
- 04/20/2012 Marco 09650 - PMT Gain Final graph of the PMT #09650 gain.
Error creating thumbnail: Unable to save thumbnail to destination
536 05/01/2012 Marco/Felipe 09650 -1.78 Prot. light tight Prototype inserted inside a dark box and signal passing through a discriminator to generate trigger. Discriminator adjusted to have about 0.5kHz rate.
Error creating thumbnail: Unable to save thumbnail to destination
537 05/01/2012 Marco/Felipe 09650 -1.78 Prot. light tight Same as 536, just turning off all the lights inside the lab to make sure there is no light leak.
Error creating thumbnail: Unable to save thumbnail to destination
538 05/01/2012 Marco/Felipe 09650 -1.78 Prot. light tight A small piece of scintillator material (3cm x 7cm) placed inside the prototype, to see if something change.
Error creating thumbnail: Unable to save thumbnail to destination
539 05/01/2012 Marco/Felipe 09650 -1.78 Prot. light tight An external trigger was created to gate the ADC, making possible the pedestal identification.
Error creating thumbnail: Unable to save thumbnail to destination
544 05/08/2012 Marco 09641 -1.77 PMT Dark Current PMT changed. Inside de plastic box, analyzing the dark current for the PMT (no LED or scintillator). Discriminating signals to have rate of approx. 1.5kHz.
Error creating thumbnail: Unable to save thumbnail to destination
545 05/08/2012 Marco 09641 -1.77 PMT Dark Current Same as #544, just reducing the discriminator level to get lower signal (higher rate - approx. 3kHz).
Error creating thumbnail: Unable to save thumbnail to destination
547 05/08/2012 Marco 09641 -1.77 PMT Gain Flashing LED in front of the PMT to get the gain. Approx. 1kHz data rate.
Error creating thumbnail: Unable to save thumbnail to destination
548 05/08/2012 Marco 09641 -1.77 PMT Gain Same as #547, changing LED intensity.
Error creating thumbnail: Unable to save thumbnail to destination
549 05/08/2012 Marco 09641 -1.77 PMT Gain Delaying signal (with a long cable) to gate the ADC in a zero level, taking the pedestal position.
Error creating thumbnail: Unable to save thumbnail to destination
550 05/08/2012 Mike/Marco 09650 -1.77 Prototype noise PMT changed. Prototype inside a dark box. Signal passing through a Linear Fan-Out, then going to the Discriminator to generate the gate (for ADC) and going to ADC channel #3. Discriminator level adjusted to have counting rate of about 1kHz. Cables used to delay signal. Pedestal position from run #553 and SEP to pedestal calibration from run #479.
Error creating thumbnail: Unable to save thumbnail to destination
551 05/08/2012 Mike/Marco 09650 -1.77 Prototype noise Same as #550, but with the dark box opened. Lab light hitting the prototype.
Error creating thumbnail: Unable to save thumbnail to destination
552 05/08/2012 Mike/Marco 09650 -1.77 Prototype noise Signal delayed by a long cable, so triggering in the zero level to analyze the pedestal position on the ADC.
Error creating thumbnail: Unable to save thumbnail to destination
556 05/14/2012 Nathaniel 09561 -1.699 PMT Gain LED Intensity @ 2.6V, Duration 15 ns, data rate 1 kHz.
Error creating thumbnail: Unable to save thumbnail to destination
558 05/14/2012 Nathaniel 09561 -1.798 PMT Gain LED Intensity @ 2.6V, Duration 15 ns, data rate 1 kHz.
Error creating thumbnail: Unable to save thumbnail to destination
559 05/14/2012 Nathaniel 09561 -1.901 PMT Gain LED Intensity @ 2.6V, Duration 15 ns, data rate 1 kHz.
Error creating thumbnail: Unable to save thumbnail to destination
560 05/14/2012 Nathaniel 09561 -2.003 PMT Gain LED Intensity @ 2.6V, Duration 15 ns, data rate 1 kHz.
Error creating thumbnail: Unable to save thumbnail to destination
561 05/14/2012 Nathaniel 09561 -2.101 PMT Gain LED Intensity @ 2.6V, Duration 15 ns, data rate 1 kHz.
Error creating thumbnail: Unable to save thumbnail to destination
562 05/14/2012 Nathaniel 09561 -2.199 PMT Gain LED Intensity @ 2.6V, Duration 15 ns, data rate 1 kHz.
Error creating thumbnail: Unable to save thumbnail to destination
563 05/14/2012 Nathaniel 09561 -2.301 PMT Gain LED Intensity @ 2.6V, Duration 15 ns, data rate 1 kHz.
Error creating thumbnail: Unable to save thumbnail to destination
556-563 05/14/2012 Nathaniel 09561 Various PMT Gain Here is the Log Scale Gain Plot for PMT 9561 (Low Gain) The Table of values may be accessed at File:PMT Gain 1.xls File:PMT Gain Plot.gif
566 05/15/2012 Marco 09561 -2.000 PMT Scanning LED Intensity @ 5.5V, Duration 20 ns. It is a low gain PMT, according to Yerevan classification. Using mu-metal shielding. Reference PMT unplugged. Using File: PMT_test_100x100_1k_v2.txt. There is a small perturbation on the position of the pixels, due to the zigzag programming. This perturbation is due to the position of the end-of-course switch at the end of the rail, it is a little bit after the last scanned position. To correct that, the code is moving the motor to achieve this end before starting the new line.
Error creating thumbnail: Unable to save thumbnail to destination
567 05/15/2012 Marco 09561 -1.803 PMT Scanning Same as 566, just changing high voltage.
Error creating thumbnail: Unable to save thumbnail to destination
568 05/15/2012 Marco 09561 -2.200 PMT Scanning Same as 566, just changing high voltage.
Error creating thumbnail: Unable to save thumbnail to destination
571-583 05/15/2012 Nathaniel [No PMT] [No HV] Attenuation Factor Comparing Measurements from Ocilloscope and ADC. Ran into difficulty. There is instead Amplification! File:Att fac darkbox1.xls upper portion of spreadsheet
584-595 05/21/2012 Nathaniel/Marco [No PMT] [No HV] Correction/Calibration In order to address the above difficulty, we widen the signal to exceed the size of the gate. Then we ran these tests in which we measure the same signal in Oscilloscope and ADC to find a correction factor Same as above, see lowest portion of spreadsheet
585-610 05/21/2012 Nathaniel/Marco [No PMT] [No HV] Attenuation Factor We use the newly found calibration factor and redesigned test to find a correct attenuation factor Same as above, see middle portion of spreadsheet.
613 05/22/2012 Marco 09561 -1.796 PMT Scanning As the zigzag in the scanning was causing a small perturbation on the position definition of the pixels (see run #566), the code for the step-motors was changed. The new code (PMT_test_100x100_1k_v3_return.txt) scan all the lines in the same direction, coming back to the zero position before starting each line scanning. There is no more zigzag. LED Intensity @ 5.5V, Duration 20 ns. It is a low gain PMT, according to Yerevan classification. Using mu-metal shielding. Reference PMT in channel 2. Signal in number of photoelectrons; using SEP-PED (in number of channels) from run #558 as the conversion factor. Reference PMT signal: Ref_613 - not used for analysis.
Error creating thumbnail: Unable to save thumbnail to destination
614 05/23/2012 Marco 09561 -2.000 PMT Scanning Same as 613, just changing high voltage. To convert it to number of photoelectrons, using SEP-PED (in number of channels) from run #560. Reference PMT signal: Ref_614.
Error creating thumbnail: Unable to save thumbnail to destination
615 05/23/2012 Marco 09561 -2.200 PMT Scanning Same as 613, just changing high voltage. To convert it to number of photoelectrons, using SEP-PED (in number of channels) from run #560. Reference PMT signal (NOT GOOD!): Ref_615.
Error creating thumbnail: Unable to save thumbnail to destination
627 05/24/2012 Marco 09561 -2.000 PMT Scanning To have more photoelectrons, three (of four) attenuation paper (pieces of office paper placed inside the LED case, before the colimator) were removed. Just one attenuator (paper) was keep. The pulse high was decreased to 8 (in the pot) and 3.5 in time (also in the pot). Scan matrix: 100x100x1k. As the LED was disconnected to change the number of collimator papers, the re-connection didn't work and a bad contact occurred. Bad run. -
631 05/24/2012 Marco 09561 -1.800 PMT Scanning Same as run #627, fixing the bad connection of the LED terminals. LED with pot 8 (full pot 5V) @ pot 4 (full pot 15 ns). Tried to fix bad contact in LED. Reference PMT signal: Ref_631. Still have problems with LED.
Error creating thumbnail: Unable to save thumbnail to destination